OpenStudy (anonymous):

lim (x^2(siny)^2)/((7x^2)+(y^2)) (x,y) goes to (0,0)

OpenStudy (anonymous):

DNE d/dx=0 d/dy=\[\infty\]

OpenStudy (turingtest):

no I think the right approach is to take the path to (0,0) along y=x http://tutorial.math.lamar.edu/Classes/CalcIII/Limits.aspx